OpenStudy (anonymous):

In january each of 60 ppl purchased a $500 washing machine. in february 10% fewer ppl purchased the same washing machine ahtat increased in price by 20%. what is the change in sales from Jan to Feb?

OpenStudy (anonymous):

I do not understand this question.

OpenStudy (anonymous):

(Jan.)60 people purchased a $ 500 washing machine (500 x 60) = 30000 (feb)10% fewer people purchased the same washing machine (.10)(60) = 6. And 60 - 6 = 54 people. Increased in price by 20% -- >(500)(.20) = 100. And 100 + 500 = 600. So, 54 people purchased the washing machine at $ 600 = (54)(600) = $ 32,400. <--sales for Feb There was a (32,400 - 30,000) = $2,400 change(increase) in sales. Is this the answer you want or do you want the percent increase ?